# LODGE: Level-of-Detail Large-Scale Gaussian Splatting with Efficient Rendering

### v1

@@ A.2 Implementational Details
LOD structure optimization steps. In this section we detail the exact steps and parameters used for creating the LOD gaussians. Given a sequence of depth thresholds $d_{l}$, we build the LOD sets $\mathcal{G}^{(l)}$ from the finest to the coarsest level. We start from the set $\mathcal{G}^{(0)}$, iteratively adding more levels. At step $l$, we construct the set $\mathcal{G}^{(l)}$ via applying a 3D smoothing filter to all Gaussians in $\mathcal{G}^{(0)}$. Note that the smoothing filter is only added to the Gaussian function, whereas the parameters remain unchanged. Therefore, the optimization process cannot decrease Gaussians’ sizes below the 3D filter size. Next, we prune all Gaussians in $\mathcal{G}^{(l)}$ using the importance score pruning and then perform 1 00010001\,0001 000 optimization steps. The optimization steps use the LOD rendering procedure (with LOD levels up to $l$) described in the previous section, with the modification that depth $d_{l}$ is replaced by a random number drawn from a uniform distribution $\mathcal{U}(0.7d_{l},1.3d_{l})$. This makes the representation more robust towards cameras lying outside of the training trajectory. We use the same loss function (DSSIM+L1) as during the standard optimization, however, we only update parameters of the Gaussians in set $G_{l}$. This process of pruning and fine-tuning is repeated in total three times, with importance score thresholds of $0.2\gamma$, $0.6\gamma$, and $\gamma$, where $\gamma$ is a hyperparameter. Therefore, we need overall $N_{\text{levels}}\cdot 3\cdot 1\,000$ optimization iterations.
@@
[…]
where max_radii_2D is the largest radius the Gaussian projects into (since the last densification). Moreover, instead of averaging the gradient statistics, we take the maximum. We use a gradient threshold of $0.015$ for the Hierarchical 3DGS [10] dataset and $0.03$ for the ZipNeRF [1] dataset. All models were trained for for 30 0003000030\,00030 000 iterations. We reset opacity every 3 00030003\,0003 000 steps until iteration 15 0001500015\,00015 000 and apply densification every $300$ steps from step $600$ to 15 0001500015\,00015 000. Importance score pruning is performed at steps 8 00080008\,0008 000, 16 0001600016\,00016 000, and 24 0002400024\,00024 000 with threshold of $0.02$ for all scenes, except Campus, where we use lower threshold of $0.01$.
@@ A.3 Datasets
Hierarchical 3DGS dataset [10].. The public release of the dataset contains two urban scenes (SmallCity and Campus). To collect the dataset, authors used a bicycle helmet on which they mounted 6 GoPro HERO6 Black cameras (5 for the Campus scene). The SmallCity scene was collected by riding a bicycle at around $6-7$km/h, while Campus was captured on foot wearing the helmet. Poses were estimated using COLMAP with custom parameters and hierarchical mapper. Additinal per-chunk bundle adjustment was performed, see Hierarchical 3DGS [10]. In our experiments, we use the official train/test split provided by the authors. We also use the official segmentation masks provided with the dataset to remove license plates, pedestrians, and moving cars.
@@ A.4 Evaluation Protocol & Baselines
We report standard PSNR, SSIM, and LPIPS metrics. To ensure a fair evaluation and reproducibility, before computing metrics, we round the predictions to uint8 range. For the Hierarchical 3DGS dataset, we mask the prediction and the ground truth image by the provided mask - replacing the pixels inside the mask with black color. For SSIM, we use commonly used default values and average across pixels and color channels. For LPIPS [39], we use the VGG network (unlike some baselines which used AlexNet with lower values). To compute FPS, we compute rendering times for each testing image. We compute each per-image rendering time $7$ times and take a median to ensure precise measurement. The resulting FPS is then one over average rendering time for all test images. In the rendering time, we only include the rendering part, not the time required to move the data from GPU memory at the end of the rendering. When reporting the number of Gaussians #$\#G$, we report the average number of Gaussian loaded in GPU memory when rendering. For methods which constructs LOD splits on the fly this is the total number of Gaussians in all LOD sets (if applicable). Note, that for Scaffold-GS [19] and Octree-GS [27], some attributes (colors) are not stored explicitly, but are decoded on-the-fly lowering memory requirements. All baselines and our method were trained and evaluated on a single NVIDIA A100 SXM4 40 GB GPU.
@@
Baselines. For the baselines, we include the following: Zip-NeRF [1] which is the current state-of-the-art on the Zip-NeRF dataset. Note, that this method was trained for 200K iterations - much longer than other baselines and does not achieve real-time rendering. We only include it for reference. RadSplat [24] uses Zip-NeRF for initialization and is, therefore, much more expensive to train than other baselines. It uses the same strategy as our method to remove less important Gaussians during training. The method used Zip-NeRF checkpoint trained for 200K and performed additional 45K training iterations to optimize 3DGS representation. 3DGS [9] is the basis for most other methods. We use the default parameters, training for 30K iterations. Mip-Splatting [35] augments 3DGS with antialiasing 2D and 3D filters. In our method, we also use the same 2D filter and we use the 3D filter when building LOD levels. Scaffold-GS [19] reduces memory requirements by decoding attributes of Gaussians on the fly from MLP. Therefore it is important baseline for rendering speed as our method stores all attributes explicitly. H3DGS [10] starts from a 3DGS representation optimized at the finest level (same as ours) and builds a representation by iteratively merging Gaussians, however, at render time all Gaussians need to be kept in memory and LOD split needs to be built with every rendering call. The method is currently SoTA for the Hierarchical 3DGS dataset. The FLOD [28], and Octree-GS [27] are both recent LOD 3DGS-based approaches. Octree-GS was trained for 40K iterations, while FLOD used 100K iterations for the 5 LOD levels. At rendering, we use the selective 3, 4, 5 rendering split. For all baselines we modified the code to accept masks (masking loss gradient propagation in masked regions). We have performed a limited hyperparameter search to get best performance for the datasets.

### v2

@@ A.2 Implementational Details
LOD structure optimization steps. In this section we detail the exact steps and parameters used for creating the LOD gaussians. Given a sequence of depth thresholds $d_{l}$, we build the LOD sets $\mathcal{G}^{(l)}$ from the finest to the coarsest level. We start from the set $\mathcal{G}^{(0)}$, iteratively adding more levels. At step $l$, we construct the set $\mathcal{G}^{(l)}$ via applying a 3D smoothing filter to all Gaussians in $\mathcal{G}^{(0)}$. Note that the smoothing filter is only added to the Gaussian function, whereas the parameters remain unchanged. Therefore, the optimization process cannot decrease Gaussians’ sizes below the 3D filter size. Next, we prune all Gaussians in $\mathcal{G}^{(l)}$ using the importance score pruning and then perform 1 0001\,000 optimization steps. The optimization steps use the LOD rendering procedure (with LOD levels up to $l$) described in the previous section, with the modification that depth $d_{l}$ is replaced by a random number drawn from a uniform distribution $\mathcal{U}(0.7d_{l},1.3d_{l})$. This makes the representation more robust towards cameras lying outside of the training trajectory. We use the same loss function (DSSIM+L1) as during the standard optimization, however, we only update parameters of the Gaussians in set $G_{l}$. This process of pruning and fine-tuning is repeated in total three times, with importance score thresholds of $0.2\gamma$, $0.6\gamma$, and $\gamma$, where $\gamma$ is a hyperparameter. Therefore, we need overall $N_{\text{levels}}\cdot 3\cdot 1\,000$ optimization iterations.
@@
$\text{grad\_position\_2D}\cdot\text{max\_radii\_2D}\cdot(\text{opacity})^{1/5}>=\text{threshold}\,$, (6)
where max_radii_2D is the largest radius the Gaussian projects into (since the last densification). Moreover, instead of averaging the gradient statistics, we take the maximum. We use a gradient threshold of $0.015$ for the Hierarchical 3DGS [11] dataset and $0.03$ for the ZipNeRF [2] dataset. All models were trained for for 30 00030\,000 iterations. We reset opacity every 3 0003\,000 steps until iteration 15 00015\,000 and apply densification every $300$ steps from step $600$ to 15 00015\,000. Importance score pruning is performed at steps 8 0008\,000, 16 00016\,000, and 24 00024\,000 with threshold of $0.02$ for all scenes, except Campus, where we use lower threshold of $0.01$.
@@ A.3 Datasets
Hierarchical 3DGS dataset [11].. The public release of the dataset contains two urban scenes (SmallCity and Campus). To collect the dataset, authors used a bicycle helmet on which they mounted 6 GoPro HERO6 Black cameras (5 for the Campus scene). The SmallCity scene was collected by riding a bicycle at around 676-7km/h, while Campus was captured on foot wearing the helmet. Poses were estimated using COLMAP with custom parameters and hierarchical mapper. Additinal per-chunk bundle adjustment was performed, see Hierarchical 3DGS [11]. In our experiments, we use the official train/test split provided by the authors. We also use the official segmentation masks provided with the dataset to remove license plates, pedestrians, and moving cars.
@@ A.4 Evaluation Protocol & Baselines
We report standard PSNR, SSIM, and LPIPS metrics. To ensure a fair evaluation and reproducibility, before computing metrics, we round the predictions to uint8 range. For the Hierarchical 3DGS dataset, we mask the prediction and the ground truth image by the provided mask - replacing the pixels inside the mask with black color. For SSIM, we use commonly used default values and average across pixels and color channels. For LPIPS [41], we use the VGG network (unlike some baselines which used AlexNet with lower values). To compute FPS, we compute rendering times for each testing image. We compute each per-image rendering time $7$ times and take a median to ensure precise measurement. The resulting FPS is then one over average rendering time for all test images. In the rendering time, we only include the rendering part, not the time required to move the data from GPU memory at the end of the rendering. When reporting the number of Gaussians #G\#G, we report the average number of Gaussian loaded in GPU memory when rendering. For methods which constructs LOD splits on the fly this is the total number of Gaussians in all LOD sets (if applicable). Note, that for Scaffold-GS [20] and Octree-GS [28], some attributes (colors) are not stored explicitly, but are decoded on-the-fly lowering memory requirements. All baselines and our method were trained and evaluated on a single NVIDIA A100 SXM4 40 GB GPU.
@@
Baselines. For the baselines, we include the following: Zip-NeRF [2] which is the current state-of-the-art on the Zip-NeRF dataset. Note, that this method was trained for 200K iterations - much longer than other baselines and does not achieve real-time rendering. We only include it for reference. RadSplat [25] uses Zip-NeRF for initialization and is, therefore, much more expensive to train than other baselines. It uses the same strategy as our method to remove less important Gaussians during training. The method used Zip-NeRF checkpoint trained for 200K and performed additional 45K training iterations to optimize 3DGS representation. 3DGS [10] is the basis for most other methods. We use the default parameters, training for 30K iterations. Mip-Splatting [37] augments 3DGS with antialiasing 2D and 3D filters. In our method, we also use the same 2D filter and we use the 3D filter when building LOD levels. Scaffold-GS [20] reduces memory requirements by decoding attributes of Gaussians on the fly from MLP. Therefore it is important baseline for rendering speed as our method stores all attributes explicitly. H3DGS [11] starts from a 3DGS representation optimized at the finest level (same as ours) and builds a representation by iteratively merging Gaussians, however, at render time all Gaussians need to be kept in memory and LOD split needs to be built with every rendering call. The method is currently SoTA for the Hierarchical 3DGS dataset. The FLOD [29], CityGS [18], and Octree-GS [28] are both recent LOD 3DGS-based approaches. Octree-GS was trained for 40K iterations, while FLOD used 100K iterations for the 5 LOD levels. At rendering, we use the selective 3, 4, 5 rendering split. For all baselines we modified the code to accept masks (masking loss gradient propagation in masked regions). We have performed a limited hyperparameter search to get best performance for the datasets.